OpenStudy (anonymous):

Help!!!! Part A: What is the y-intercept of the function and what does this tell you about the car? (4 points) Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 3 and what does the average rate represent? (4 points) Part C: What would be the domain of this function if the car traveled the same rate until it reached its destination? (2 points)

OpenStudy (anonymous):

The table for it

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Please help

OpenStudy (cybershadow):

for the second part , Average rate of the change of the function is y2-y1/x2-x1 , where (x1,y1),(x2,y2) are the starting and ending points respectively therefore average rate change between x = 1 , to x = 3 wud be 750-850/3-1 = -100/2 = -50

OpenStudy (cybershadow):

Did this make any sense to u? O_O

OpenStudy (anonymous):

Wouldn't the y intercept be 900

OpenStudy (cybershadow):

Oh yes Nvm! It wont be equation but it wud be x+1 is directly proptional to y+100 , Good call there! , when x is 0 , y is 900 so y-intercept is 900

OpenStudy (anonymous):

Okay then it asks what does it tell me about the car?

OpenStudy (anonymous):

I'm not sure how to explain that

OpenStudy (cybershadow):

well it just tells u that when time is 0 , car is 900 miles away from the destination

OpenStudy (cybershadow):

Theres no other thing it could possibly tell abt the car...... .-. Dint make sense eh?

OpenStudy (anonymous):

No it makes sense and for C?

OpenStudy (cybershadow):

if car is travelling from the same rate , that means distance is going to decrease in the intervels of 50 miles until it becomes 0. 0 900 1 850 2 800 3 750 4 700 5 650 6 600 7 550 8 500 9 450 10 400 11 350 12 300 13 250 14 200 15 150 16 100 17 50 18 0 YAY! Reached the destination !, jesus christ it took so much effort to type all dat So Domain wud be all the numbers from 0 to 18 or u can write [0,18]

OpenStudy (anonymous):

Lol! Sorry thank you can you help me with more questions

OpenStudy (anonymous):

#2

OpenStudy (cybershadow):

ok well.lets call the rent y and days x then rate of change of function wud be dy/dx = 625-465/7-5 = 160/2=80 Now to write it in the standard x, y form take the dx to the otherside and integrate the equation \[\int\limits dy =80\int\limits dx \] that would make it y = 80x + c (constant)

OpenStudy (anonymous):

Okay, sorry but going back up to number one for part B what does the average rate of change represents

OpenStudy (cybershadow):

Well it represents the average speed for the first question

OpenStudy (cybershadow):

average speed of the car

OpenStudy (anonymous):

It represents the average speed of the car from one hour to the next?

OpenStudy (cybershadow):

Yeap!

OpenStudy (anonymous):

Okay back to #2

OpenStudy (cybershadow):

now going back to the second question we got y = 80 x + c now we gotta find c u can place value of x , y from the given values that is 465 and 5 that makes our equation 465 = 80 x 5 +c 465 = 400 + c c = 65 therefore our genral equation is now y = 80 x + 65

OpenStudy (cybershadow):

is that clear?

OpenStudy (anonymous):

Yes , as a function notation would I just add fx to the end of it?

OpenStudy (cybershadow):

Yes!

OpenStudy (cybershadow):

Actually in place of y , u have to write F(x)

OpenStudy (anonymous):

Yeah, okay how would I do part c

OpenStudy (cybershadow):

well just alot random values to x and see what value y will get , and mark those points in graph, for instance when x = 0 , y = 65 so our point on the graph would be (0,65) take three-four such points by assigning values to x , and join those points

OpenStudy (anonymous):

Okay what do they mean by intervals ?

OpenStudy (cybershadow):

intervals are the difference between the values of the y

OpenStudy (cybershadow):

and differnce in the values of x

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

#3

OpenStudy (cybershadow):

well slope of the function wud be dy/dx so slope of f(x) 7-(-9)/1-1 = infinite while the slope of g(x) = d/dx(3x-2) = 3

OpenStudy (anonymous):

I got 8 for the slope

OpenStudy (cybershadow):

Oh my bad 7-(-9) / 1-(-1) = 16/2 that would be 8

OpenStudy (cybershadow):

plus g(x) 's y intercept wud be 3(0)-2 = -2 and f(x)'s y intercept is -1 so F(x) has greater y intercept

OpenStudy (anonymous):

How do the slope s compare, they're both positive?

OpenStudy (cybershadow):

Yeap But g(x) has larger value of slope so it wud be less bent toward X-axis

OpenStudy (anonymous):

You mean fx has a larger slope

OpenStudy (cybershadow):

Ah yes! 8 is of F(x) right? then yes xD sorry i can hardly keep my eyes open its 8 Am here xD

OpenStudy (anonymous):

Oh, I'm sorry lol . #4

OpenStudy (cybershadow):

isnt this the same question? O_O

OpenStudy (anonymous):

Sorry

OpenStudy (cybershadow):

this relation isnt a function as a function cannot have different images for the same x in the domain

OpenStudy (anonymous):

So every x has to have their own individual output?

OpenStudy (cybershadow):

Yes! one value of x cannot have two different values of y , in the given relation 1 has two values of y (0,-1) , therefore its not a function

OpenStudy (cybershadow):

and in the second part to find f(4) just put the value of x as 4 in the equation that will give u the value of f(4)

OpenStudy (cybershadow):

and In part C an ordered pair is---> (x,y) so when x is p , y would be -p+6 so ordered pair will be (p , 6-p)

OpenStudy (cybershadow):

got it eh? O_O

OpenStudy (anonymous):

One moment I'm looking at it

OpenStudy (anonymous):

Okay I got it

OpenStudy (anonymous):

#5

OpenStudy (cybershadow):

f(w) total weight will be weight of the trucks + n x weight of each bale therefore f(w) = 4500+150n f(w) = 150(30 + n)

OpenStudy (cybershadow):

for part b put n as 15 that will make f(w) = 150(30+15) = calculate that

OpenStudy (cybershadow):

for part c , a cartoon worth 20Kg(Eh? must be daffy duck,oh waits Its carton ,idk why i read that as cartoon :/) is also added to the truck so now total wait wud be f(w) = 150(30+10) + 20 f(w) = 150(300) + 20 = 45020 (idk if thats correct, calculate) :/

OpenStudy (anonymous):

Okay, thank you so much!

OpenStudy (cybershadow):

thats all? *whew* No worriez

OpenStudy (anonymous):

You're amazing

OpenStudy (cybershadow):

gee thanks! glad to help